The area of a rectangular field is 150 yd. if it's length is 1 1/2 times its width, find the perimeter of the field?
Answer by amarjeeth123(569)

Let the width be x.
Then the length is 1.5x.
The area of a rectangular field is 150 square yd.
Area=length*width
1.5x*x=150
x^2=100
x=10 yards
Width=10 yards
Length=15 yards
Perimeter of the field=2(length+width)=2(15+10)=2*25=50 yards.